OpenStudy (anonymous):

Solve for x^2. x^4-4X^2+2=0. Please show work.

jimthompson5910 (jim_thompson5910):

Do you mean solve for x?

OpenStudy (anonymous):

No like \[x^{2}\]

OpenStudy (neo92):

take x^2 = t , then \[x^{4}-4x^{2}+2 = t^{2}-4t+2=0\] now solve the eqn \[t^{2}-4t+2\] for t

jimthompson5910 (jim_thompson5910):

well you can let z = x^2 So z^2 = x^4, which means x^4-4x^2+2=0 becomes z^2-4z+2=0 Now solve for z

OpenStudy (neo92):

@floridagirl150 got it?

OpenStudy (anonymous):

i'm still kinda confused but that helps me get started. Thanks :)